OpenStudy (anonymous):

Alright! I am getting confused now! Can someone tell me how to Estimate the percent of a number? For example. 76% of 80 = 60! I don't understand that.

OpenStudy (anonymous):

(76 / 100) * 80 = 0.76 * 80 = 60.8